The integral curves of "\frac {dx}{a} = \frac {dy}{b} = dz" is


1
Expert's answer
2021-09-21T09:51:38-0400

Let's consider "\frac{dx}{a}= \frac{dy}{b}." Antiderivatives are related by the folowing equalitie "\frac{x}{a}=\frac{y}{b}+c_1" .

Let's consider the "\frac{dx}{a}= dz" . Antiderivatives are related by the folowing equalitie "\frac{x}{a}=z+c_2."

So the answer is "\frac{x}{a}=\frac{y+c_1b}{b}=z+c_2"